cot 60 degrees
tan 210 degrees
sec 210
csc 210
cot 210
cos 11pi/6
cot 3pi/2
tan 2pi/3
csc 5pi/6

please help if anyone can do it i need it by thursday 8/23/06 by 6:30 tomorrow morning please help!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

Answer by venugopalramana(3286)

cot 60 degrees=1/SQRT(3)
tan 210 degrees=TAN(180+30)=TAN(30)=1/SQRT(3)
sec 210=SEC(180+30)=-SEC(30)=-2/SQRT(3)
csc 210=COSEC(180+30)=-COSEC(30)=-2
cot 210=COT(180+30)=COT(30)=SQRT(3)
cos 11pi/6=COS(2PI-PI/6)=COS(PI/6)=SQRT(3)/2
cot 3pi/2=COT(PI+PI/2)=COT(PI/2)=0
tan 2pi/3=TAN(PI-PI/3)=-TAN(PI/3)=-SQRT(3)
csc 5pi/6 =COSEC(PI-PI/6)=COSEC(PI/6)=2